Throw a single die and let the random variable X be the number on the die. construct the probability distribution including possible value of X and find the mean of the distribution.

1
Expert's answer
2022-03-07T13:28:09-0500

There are "6" possible outcomes. Since the events are equally likely than


"P(X=1)=P(X=2)=P(X=3)=P(X=4)"

"=P(X=5)=P(X=6)=1\/6"

Construct the probability distribution


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 p(x) & 1/6 & 1/6 & 1/6 & 1/6 & 1/6 & 1/6 
\end{array}"


"mean=E(X)=1(\dfrac{1}{6})+2(\dfrac{1}{6})+3(\dfrac{1}{6})"

"+4(\dfrac{1}{6})+5(\dfrac{1}{6})+6(\dfrac{1}{6})=3.5"